A 60 mL flask is filled with mercury at 00C. When the temperature rises to 800C, 1.47 g of mercury spills out of the flask. Assuming the volume of the flask is constant, find the density of mercury at 800C if its density at 00C is 13,645 kg/m3.
1
Expert's answer
2021-08-23T12:35:27-0400
"M=\rho V\\\rho'=\frac{M-\Delta m}{V}\\\rho'=\rho-\frac{\Delta m}{V}
\\\rho'=13645-\frac{1470}{60}=13620.5\frac{kg}{m^3}"